### v1

@@ Abstract
While Large Language Models (LLMs) have demonstrated remarkable potential in time series forecasting, their practical deployment remains constrained by excessive computational demands and memory footprints. Existing LLM-based approaches typically suffer from three critical limitations: (1) Inefficient parameter utilization in handling numerical time series patterns; (2) Modality misalignment between continuous temporal signals and discrete text embeddings; and (3) Inflexibility for real-time expert knowledge integration. We present Small but Mighty Enhancing Time Series (SMETimes), the first systematic investigation of sub-3B parameter Small Language Models (SLMs) for efficient and accurate time series forecasting. Our approach centers on three key innovations: (1) A statistically-enhanced prompting mechanism that bridges numerical time series with textual semantics through descriptive statistical features; (2) A adaptive fusion embedding architecture that aligns temporal patterns with language model token spaces through learnable parameters; And (3) a dynamic mixture-of-experts framework enabled by SLMs’ computational efficiency, adaptively combining base predictions with domain-specific models. Extensive evaluations across seven benchmark datasets (ETTh1/2, ETTm1/2, Weather, Solar, ECL) demonstrate that our 3B-parameter SLM achieves state-of-the-art performance on five primary datasets while maintaining 3.8$\times$ faster training and 5.2$\times$ lower memory consumption compared to 7B-parameter LLM baselines. Notably, the proposed model exhibits better learning capabilities, achieving 12.3% lower MSE than conventional LLM. Ablation studies validate that our statistical prompting and cross-modal fusion modules respectively contribute 15.7% and 18.2% error reduction in long-horizon forecasting tasks. By redefining the efficiency-accuracy trade-off landscape, this work establishes SLMs as viable alternatives to resource-intensive LLMs for practical time series forecasting. Code and models are available at https://github.com/xiyan1234567/SMETimes.
@@
Small Language Models, Statistically-enhanced Prompting, Adaptive Fusion Embedding
@@ 1 Introduction
Time series forecasting stands as a cornerstone of modern decision-making systems, with critical applications spanning energy management (petrucci2022modelling, ), financial markets (leonard2001promotional, ), climate modeling (schneider1974climate, ), and intelligent transportation (li2015demand, ). Traditional approaches often rely on domain-specific statistical models (box1994time, ) or deep neural networks (liu2023itransformer, ), which require substantial computational resources and extensive domain expertise. While large language models (LLMs) have recently demonstrated remarkable capabilities in time series forecasting (jin2023time, ; zhou2023ptse, ), their practical deployment faces significant challenges due to prohibitive computational costs and memory footprints (touvron2023llama, ).
Figure 1: Performance-efficiency Trade-off Comparison on ETTh1 (zhou2021informer, ) Dataset. Our SLM variants (blue) achieve competitive MSE with significantly lower training time and memory footprint compared to conventional LLM-based approaches. Bubble size represents relative memory consumption.
The recent proliferation of LLM-based forecasting methods (woo2024unified, ; gruver2023large, ) has revealed three fundamental limitations: (1) Massive parameter counts (typically $>$7B) lead to inefficient training/inference; (2) Inadequate alignment between numerical time series patterns and textual embeddings; And (3) limited flexibility for integrating domain-specific expert knowledge. As shown in Figure 1, conventional LLM approaches like Time-LLM (jin2023time, ) (22.33G) and AutoTimes liu2025autotimes (23.12G) incur substantial resource costs despite comparable MSE performance to our 1B-parameter SLM (4.33G). This efficiency gap becomes particularly critical in real-world deployment scenarios with hardware constraints.
To alleviate the computational inefficiency of large language models (LLMs) in time series forecasting, we propose a feature fusion strategy through Small Language Models (SLMs) that strategically reduce model scale while incorporating targeted architectural innovations to achieve better efficiency. Our key insight lies in three synergistic components: (1) Statistically-enhanced prompting that bridges numerical and textual domains, (2) A adaptive fusion mechanism for time series embeddings, And (3) dynamic mixture-of-experts (MoE) integration enabled by SLMs’ lightweight architecture. Extensive experiments across seven benchmark datasets (ETTh1/2 (zhou2021informer, ), ETTm1/2 (zhou2021informer, ), Weather (wu2021autoformer, ), Solar (lai2018modeling, ), ECL (wu2021autoformer, )) demonstrate that our 1B-parameter SLM achieves state-of-the-art results on five primary datasets while maintaining competitive performance on the remaining three.
@@
  • To the best of our knowledge, this is the first work to develop and evaluate a framework that applies small language models (SLMs) to time series forecasting tasks. Through targeted architectural modifications, we demonstrate that compact models achieve performance parity with 7B-parameter large language models while attaining 3.2$\times$ accelerated training convergence and 5.1$\times$ reduced memory footprint, substantially enhancing deployment feasibility in resource-constrained environments.
  • We introduce a new prompting methodology integrating temporal statistics with domain-specific contextual metadata. Empirical validation reveals this statistically-enhanced approach reduces mean squared error by 12.7% compared to conventional textual prompting baselines through systematic ablation analysis.
  • Our adaptive fusion architecture addresses the intrinsic modality gap between continuous time series embeddings and discrete token representations. By implementing learnable projective transformations coupled with attention-based feature alignment, the proposed mechanism yields 9.3% accuracy improvement on extended forecasting horizons compared to standard embedding approaches.
  • The developed dynamic mixture-of-experts framework synergistically combines base model predictions with established temporal modeling techniques, including ARIMA (shumway2017arima, ) and Prophet (chen2009overview, ). This hybrid architecture achieves 4.8% MSE reduction while maintaining 2.1$\times$ faster inference speeds compared to monolithic LLM implementations, demonstrating effective balance between computational efficiency and forecasting precision.
The remainder of this paper is organized as follows: Section 2 reviews relevant work in LLM-based forecasting and model compression. Section 3 details our SLM architecture and technical innovations. Sections 4 present comprehensive experiments and ablation studies. We conclude with discussions of societal impacts and future directions in Section 7.
@@ 2.1.1 Traditional Approaches
Traditional time series forecasting has long relied on domain-specific statistical models and classical machine learning techniques. Methods such as ARIMA (shumway2017arima, ) and its variants (e.g., SARIMA (dubey2021study, )) leverage autoregressive and moving average components to model temporal dependencies but struggle with nonlinear patterns and multivariate data (box1994time, ). Exponential smoothing (winters1960forecasting, ) and state space models (SSMs) (liu2023koopa, ) further incorporate trend and seasonality decomposition, yet their rigidity limits adaptability to complex real-world scenarios. With the rise of deep learning, architectures like LSTMs (hochreiter1997long, ) and Temporal Convolutional Networks (TCNs) (bai2018empirical, ) emerged as powerful tools for sequence modeling. Transformers (vaswani2017attention, ), initially designed for NLP, were later adapted for time series (e.g., Informer (zhou2021informer, ), Autoformer (wu2021autoformer, )) to capture long-range dependencies via self-attention. While these methods achieve strong performance on narrow tasks, they require extensive domain expertise, task-specific tuning, and large-scale training data, limiting their generalizability across diverse applications (wen2023transformers, ).
2.1.2 LLM-Based Approaches
@@
Recent advances in large language models (LLMs) have inspired their adaptation to time series forecasting. Pioneering works like FPT (zhou2023one, ) and Time-LLM (jin2023time, ) demonstrated that LLMs pretrained on textual data can be repurposed for temporal modeling through cross-modal alignment. For instance, LLMTime (gruver2023large, ) treats time series as numerical tokens, enabling zero-shot forecasting via LLMs’ inherent pattern recognition capabilities. Methods such as PromptCast (xue2023promptcast, ) and TEMPO (cao2023tempo, ) further refine prompting strategies to bridge numerical and textual modalities. However, these approaches inherit critical limitations from their reliance on massive LLMs (e.g., GPT-3 (floridi2020gpt, ), Llama (touvron2023llama, )): (1) Excessive computational costs (e.g., AutoTimes (liu2025autotimes, ) requires 23.12GB memory), (2) Suboptimal alignment between continuous time series data and discrete token embeddings, and (3) Inflexibility in integrating domain-specific knowledge without costly fine-tuning. While autoregressive LLM-based methods like Time-LLM (jin2023time, ) achieve variable-length predictions, they suffer from quadratic attention complexity and high inference latency, rendering them impractical for resource-constrained environments.
2.2 Small Language Models for Time Series
@@
Recent advancements in SLM-based time series analysis have primarily addressed classification and edge deployment challenges, yet critical gaps persist in forecasting tasks. While Voice2Series (yang2021voice2series, ) and EdgeTS (chen2023edge, ) demonstrate the feasibility of parameter-efficient SLMs for temporal pattern recognition, their focus on short-term classification or latency reduction overlooks the intrinsic complexities of multi-step forecasting. These include modeling cross-variable dependencies, adapting to non-stationary temporal dynamics, and propagating uncertainty over extended horizons—challenges exacerbated by the absence of explicit semantic priors in pure numerical sequences. Existing LLM-based forecasting frameworks (jin2023time, ; xue2023promptcast, ) partially address these issues through language-aligned prompting but introduce modality misalignment when fusing time series tokens with textual instructions. Our work bridges this gap by leveraging SLMs’ architectural flexibility to natively encode temporal semantics through quantized embeddings and timestamp-informed attention mechanisms. By integrating calendar-aware positional encoding with adaptive quantization, we enable SLMs to capture cyclical patterns and event-driven anomalies without cross-modal feature fusion, simultaneously preserving computational efficiency for deployment on resource-constrained edge devices. This approach extends the SLM paradigm beyond classification-centric designs, addressing the understudied trade-off between long-term dependency modeling and real-time inference in forecasting applications.
@@ 3 Methodology
As shown in Figure 2, the proposed SMETimes framework features three principal components: (1) Statistically-enhanced prompt engineering, (2) Adaptive fusion with dynamic gating mechanisms, and (3) Dynamic Mixture-of-Experts (MoE) prediction head. The methodology systematically elaborates these components through dedicated technical sections: Section 3.1 formalizes the numerical-textual alignment process via statistical prompting, Section 3.2 specifies the implementation details of the dynamic gating architecture, and Section 3.3 delineates the parameter allocation strategy for the MoE specialization module. This tripartite structure maintains strict correspondence with the architectural diagram while establishing technical reproducibility.
Figure 2: Architecture of the proposed SMETimes framework, featuring three core innovations: (1) Statistically-enhanced prompt engineering for numerical-textual alignment, (2) Adaptive fusion with dynamic gating mechanisms, And (3) dynamic Mixture-of-Experts (MoE) prediction head for efficient specialization.
3.1 Statistically-enhanced Prompt
To bridge the gap between numerical time series patterns and natural language semantics, we propose a statistically-enhanced descriptor mechanism that generates linguistically interpretable embeddings through domain-specific statistical features and timestamp contextualization. Given a univariate time series $\mathbf{X}\!=\!\{x_{1},\dots,x_{T}\}\in\mathbb{R}^{T}$, we first partition it into $N$ non-overlapping segments following the sliding window strategy in liu2023itransformer :
@@ 3.1 Statistically-enhanced Prompt
where $S$ denotes the segment length controlling temporal granularity. The selection of $S$ follows domain-specific periodicity validated through sensitivity analysis in Section 4. And $N$ represents the total segment count.
Each segment $\mathbf{s}_{i}$ undergoes parallel processing through two complementary descriptors: temporal contextualization and statistical characterization. The timestamp descriptor $\mathcal{T}_{i}$ converts the start/end timestamps of $\mathbf{s}_{i}$ into natural language phrases (e.g., “03-Jan-2023 08:00 to 03-Jan-2023 12:00”) via $\operatorname{TimeStampDescriptor}(\cdot)$, while the statistical descriptor $\mathcal{S}_{i}$ extracts distributional properties using $\operatorname{StatDescriptor}(\cdot)$:
$\mathcal{S}_{i}=\operatorname{StatDescriptor}(\mathbf{s}_{i})=[\mu(\mathbf{s}_{i}),\sigma(\mathbf{s}_{i}),\nabla(\mathbf{s}_{i})]\in\mathbb{R}^{3}$ (2)
@@
where $\mu(\cdot)$, $\sigma(\cdot)$, and ($\nabla$($\cdot$)∇ ( ⋅ ) compute the mean, standard deviation, and linear trend coefficient respectively, inspired by feature engineering in nie2022time . The concatenated descriptor $\mathcal{T}_{i}\oplus\mathcal{S}_{i}$ is encoded through frozen LLM layers from pre-trained models vaswani2017attention , with $\operatorname{SelectFinal}(\cdot)$ extracting the final-layer representation $\mathbf{TE}_{i}\in\mathbb{R}^{D}$ ($D$: LLM hidden dimension). This hybrid design ensures $\mathbf{TE}_{i}$ encodes both numerical regularity and contextual temporality, establishing cross-modal alignment while avoiding redundant LLM computations during training through offline pre-computation jin2023time . The frozen LLM parameters preserve linguistic priors while the trainable $\operatorname{SelectFinal}(\cdot)$ adapts embeddings to time series forecasting objectives.
3.2 Adaptive Fusion
To harmonize numerical segment embeddings $\mathbf{SE}_{i}$ and textual descriptors $\mathbf{TE}_{i}$, we propose a learnable fusion layer that adaptively adjusts modality contributions. Let $\mathbf{SE}_{i}=\operatorname{SegmentEmbedding}(\mathbf{s}_{i})\in\mathbb{R}^{D}$ encode local temporal patterns through convolutional filters nie2022time , while $\mathbf{TE}_{i}\in\mathbb{R}^{D}$ from Section 3.1 captures global semantic contexts. Inspired by dynamic feature fusion in multimodal learning jin2023time , we introduce trainable parameter $\theta\in\mathbb{R}$ with sigmoid activation:
$\mathbf{E}_{i}=\alpha\mathbf{SE}_{i}+(1-\alpha)\mathbf{TE}_{i},\ \ \alpha=\sigma(\theta)$ (3)
where $\sigma(\cdot)$ denotes the sigmoid function constraining $\alpha\in[0,1]$. This enables instance-dependent weighting - segments with stable patterns lean towards $\mathbf{SE}_{i}$ ($\alpha\to 1$), while volatile series emphasize $\mathbf{TE}_{i}$ ($\alpha\to 0$). For batch $B\subset\{1,...,N-1\}$, fused embeddings $\{\mathbf{E}_{i}\}_{i\in B}$ are processed through frozen LLM layers vaswani2017attention :
$\{\hat{\mathbf{E}}_{2},...,\hat{\mathbf{E}}_{B+1}\}=\operatorname{LLMLayers}(\{\mathbf{E}_{i}\}_{i\in B})$ (4)
preserving temporal causality through autoregressive attention. Compared to static concatenation in liu2024unitime , our single-parameter design achieves better interpretability with negligible computation overhead.
3.3 Dynamic Mixture-of-Experts
To model diverse temporal patterns, we implement sparse gated MoE architecture shazeer2017outrageously with $K$ domain experts. Each batch element $\hat{\mathbf{E}}_{i}\in\mathbb{R}^{D}$ generates gating weights through:

### v2

@@ Abstract
While Large Language Models (LLMs) have demonstrated remarkable potential in time series forecasting, their practical deployment remains constrained by excessive computational demands and memory footprints. Existing LLM-based methods typically suffer from three critical limitations: (1) Inefficient parameter utilization in handling numerical time series patterns; (2) Modality misalignment between continuous temporal signals and discrete text embeddings; and (3) Inflexibility for real-time expert knowledge integration. We present Small but Mighty Enhancing Time Series (SMETimes), the first systematic investigation of Small Language Models with sub-3B parameters (SLM) for efficient and accurate time series forecasting. Our method centers on three key innovations: (1) A statistically enhanced prompt structure that bridges numerical time series with textual semantics through descriptive statistical features; (2) An adaptive fusion embedding structure that aligns temporal patterns with language model token spaces through learnable parameters; And (3) a dynamic mixture-of-experts structure enabled by SLMs’ computational efficiency, adaptively combining base predictions with domain-specific models. Extensive evaluations across seven benchmark datasets (ETTh1/2, ETTm1/2, Weather, Solar, ECL) demonstrate that our 3B-parameter SLM achieves state-of-the-art performance on five primary datasets while maintaining 3.8$\times$ faster training and 5.2$\times$ lower memory consumption compared to 7B-parameter LLM baselines. In particular, the proposed model exhibits better learning capabilities, achieving 12.3% lower MSE than conventional LLM. Ablation studies validate that our statistically enhanced prompt structure and adaptive fusion embedding structure contribute, respectively, to the reduction of 15.7% and 18.2% errors in long-horizon forecasting tasks. By redefining the efficiency-accuracy trade-off landscape, this work establishes SLMs as viable alternatives to resource-intensive LLMs for practical time series forecasting. The code and models are available at https://github.com/xiyan1234567/SMETimes.
@@
Small Language Models, Statistically Enhanced Prompt, Adapti- ve Fusion Embedding, Dynamic Mixture-of-Experts
@@ 1 Introduction
Time series forecasting is the cornerstone of modern decision-making systems, with critical applications spanning energy management (petrucci2022modelling, ), financial markets (zhao2023novel, ), climate modeling (schneider1974climate, ) and intelligent transportation (li2015demand, ). Traditional methods often rely on domain-specific statistical models (box1994time, ) or deep neural networks (liu2023itransformer, ), which require substantial computational resources and extensive domain expertise. Although Large Language Models (LLMs) have recently demonstrated remarkable capabilities in time series forecasting (jin2023time, ; zhou2023ptse, ), their practical deployment faces significant challenges due to prohibitive computational costs and memory footprints (touvron2023llama, ).
Figure 1: Performance-efficiency Trade-off Comparison on ETTh1 (zhou2021informer, ) Dataset. Our SLM variants (blue) achieve competitive MSE with significantly lower training time and memory footprint compared to conventional LLM-based methods. Bubble size represents relative memory consumption.
The recent proliferation of LLM-based forecasting methods (woo2024unified, ; gruver2023large, ) has revealed three fundamental limitations: (1) Massive parameter counts (typically $>$7B) lead to inefficient training/inference; (2) Inadequate alignment between numerical time series patterns and textual embeddings; And (3) limited flexibility to integrate domain-specific expert knowledge. As shown in Fig. 1, conventional LLM methods such as Time-LLM (jin2023time, ) (22.33G memory) and AutoTimes liu2025autotimes (23.12G memory) incur substantial resource costs despite comparable MSE performance to our 1B-parameter SLM (4.33G memory). This efficiency gap becomes particularly critical in real-world deployment scenarios with hardware constraints.
To alleviate the computational inefficiency of Large Language Models (LLMs) in time series forecasting, we propose a feature fusion strategy through Small Language Models (SLMs) that strategically reduce model scale while incorporating targeted architectural innovations to achieve better efficiency. Our key insight lies in three synergistic components: (1) Statistically enhanced prompt structure that bridges numerical and textual domains; (2) An adaptive fusion embedding structure for time series embeddings; And (3) a dynamic mixture-of-experts structure enabled by SLMs’ lightweight structure. Extensive experiments on seven benchmark datasets (ETTh1/2 (zhou2021informer, ), ETTm1/2 (zhou2021informer, ), Weather (wu2021autoformer, ), Solar (lai2018modeling, ), ECL (wu2021autoformer, )) demonstrate that our 1B parameter SLM achieves state-of-the-art results on five primary datasets while maintaining competitive performance on the remaining three.
@@
  • To the best of our knowledge, this is the first work to develop and evaluate a framework that applies Small Language Models (SLMs) to time series forecasting tasks. Through targeted architectural modifications, we demonstrate that compact models achieve performance parity with 7B-parameter Large Language Models (LLMs) while attaining 3.2$\times$ accelerated training convergence and 5.1$\times$ reduced memory footprint, substantially improving deployment feasibility in resource-constrained environments.
  • We introduce a new prompting methodology that integrates temporal statistics with domain-specific contextual metadata. Empirical validation reveals that this statistically enhanced prompt structure reduces the mean squared error by 12.7% compared to conventional textual prompting baselines through systematic ablation analysis.
  • Our adaptive fusion embedding structure addresses the intrinsic modality gap between continuous time series embeddings and discrete token representations. By implementing learnable projective transformations coupled with attention-based feature alignment, the proposed structure yields 9.3% accuracy improvement on extended forecasting horizons compared to standard embedding methods.
  • The dynamic mixture-of-experts structure developed synergistically combines the predictions of the base model with established temporal modeling techniques, including ARIMA (shumway2017arima, ) and Prophet (chen2009overview, ). This hybrid structure achieves a reduction of 4.8% MSE while maintaining 2.1$\times$ faster inference speeds compared to monolithic LLM implementations, demonstrating an effective balance between computational efficiency and forecast precision.
The remainder of this paper is organized as follows. Section 2 reviews traditional approaches for time series forecasting, LLM-based methodologies, and SLM analysis in this domain. Section 3 details our framework design details and technical innovations. Section 4 introduces datasets, implementation specifics, comprehensive experiments, and ablation studies, while Section 5 discusses the limitations of our model. We conclude with future research directions in Section 7.
@@ 2.1.1 Traditional Methods
Traditional time series forecasting has long relied on domain-specific statistical models and classical machine learning techniques. Methods such as ARIMA (shumway2017arima, ) and its variants (e.g. SARIMA (dubey2021study, )) leverage autoregressive and moving average components to model temporal dependencies but struggle with nonlinear patterns and multivariate data (box1994time, ). Exponential smoothing (winters1960forecasting, ) and state-space models (SSMs) (liu2023koopa, ) further incorporate trend and seasonality decomposition, yet their rigidity limits adaptability to complex real-world scenarios. With the rise of deep learning, structures such as LSTMs (hochreiter1997long, ) and Temporal Convolutional Networks (TCNs) (bai2018empirical, ) emerged as powerful tools for sequence modeling. Transformers (vaswani2017attention, ), initially designed for NLP, were later adapted for time series (cao2024ad, ; alioghli2025enhancing, ) to capture long-range dependencies through self-attention. Although these methods achieve strong performance on narrow tasks, they require extensive domain expertise, task-specific tuning, and large-scale training data, limiting their generalizability across diverse applications (wen2023transformers, ).
2.1.2 LLM-Based Methods
@@
Recent advances in Large Language Models (LLMs) have inspired their adaptation to time series forecasting. Pioneering work like FPT (zhou2023one, ) and Time-LLM (jin2023time, ) demonstrated that LLMs trained on textual data can be repurposed for temporal modeling through cross-modal alignment. For example, LLMTime (gruver2023large, ) treats time series as numerical tokens, enabling zero-shot forecasting via LLMs’ inherent pattern recognition capabilities. Methods such as PromptCast (xue2023promptcast, ) and TEMPO (cao2023tempo, ) further refine prompting strategies to bridge numerical and textual modalities. However, these methods inherit critical limitations from their reliance on massive LLMs (e.g., GPT-3 (floridi2020gpt, ), Llama (touvron2023llama, )): (1) Excessive computational costs (e.g., AutoTimes (liu2025autotimes, ) requires 23.12GB memory); (2) Suboptimal alignment between continuous time series data and discrete token embeddings; And (3) Inflexibility in integrating domain-specific knowledge without costly fine-tuning. Although autoregressive LLM-based methods like Time-LLM (jin2023time, ) achieve variable-length predictions, they suffer from quadratic attention complexity and high inference latency, rendering them impractical for resource-constrained environments.
2.2 Small Language Models for Time Series Forecasting
@@
Recent advances in SLM-based time series analysis have primarily addressed classification and edge deployment challenges, yet critical gaps persist in forecasting tasks. While Voice2Series (yang2021voice2series, ) and EdgeTS (chen2023edge, ) demonstrate the feasibility of parameter-efficient SLMs for temporal pattern recognition, their focus on short-term classification or latency reduction overlooks the intrinsic complexities of multistep forecasting. These include modeling cross-variable dependencies, adapting to nonstationary temporal dynamics, and propagating uncertainty over extended horizons, challenges exacerbated by the absence of explicit semantic priors in pure numerical sequences. The existing LLM-based forecasting structure (jin2023time, ; xue2023promptcast, ) partially addresses these issues through language-aligned prompting but introduces modality misalignment when fusing time series tokens with textual instructions. Our work bridges this gap by leveraging SLMs’ architectural flexibility to natively encode temporal semantics through quantized embeddings and timestamp-informed attention structures. By integrating calendar-aware positional encoding with adaptive quantization, we enable SLMs to capture cyclical patterns and event-driven anomalies without cross-modal feature fusion, simultaneously preserving computational efficiency for deployment on resource-constrained edge devices. This method extends the SLM paradigm beyond classification-centric designs, addressing the understudied trade-off between long-term dependency modeling and real-time inference in forecasting applications.
@@ 3 Methodology
As shown in Fig. 2, the proposed SMETimes framework features three principal components: (1) Statistically enhanced prompt structure; (2) Adaptive fusion embedding structure with dynamic gating structures; And (3) dynamic mixture-of-experts structure, the methodology systematically elaborates these components through dedicated technical sections: Section 3.1 formalizes the numerical-textual alignment process by statistical prompting, Section 3.2 specifies the implementation details of the dynamic gating framework, and Section 3.3 delineates the parameter allocation strategy for the MoE specialization module. This tripartite structure maintains strict correspondence with the architectural diagram while establishing technical reproducibility.
Figure 2: Structure of the proposed SMETimes framework, featuring three core innovations: (1) Statistically enhanced prompt structure for numerical-textual alignment; (2) Adaptive fusion embedding structure with dynamic gating structures; And (3) dynamic Mixture-of-Experts structure for efficient specialization.
3.1 Statistically Enhanced Prompt Structure
To bridge the gap between numerical time series patterns and natural language semantics, we propose a statistically enhanced prompt structure that generates linguistically interpretable embeddings through domain-specific statistical features and timestamp contextualization. Given a univariate time series $\mathbf{X}=\{x_{1},\dots,x_{T}\}\in\mathbb{R}^{T}$, we first partition it into $N$ non-overlapping segments following the sliding window strategy in this work liu2023itransformer :
@@ 3.1 Statistically Enhanced Prompt Structure
where $S$ denotes the segment length controlling temporal granularity. The selection of $S$ follows the domain-specific periodicity validated by the sensitivity analysis in Section 4.5. And $N$ represents the total segment count.
Each segment $\mathbf{s}_{i}$ undergoes parallel processing through two complementary descriptors: the timestamp descriptor and the statistical descriptor.
3.1.1 Timestamp Descriptor
To obtain the input of the timestamp descriptor and the statistical descriptor, we give a variable $\mathbf{X^{\prime}}=\{x^{\prime}_{1},\dots,x^{\prime}_{N}\}\in\mathbb{R}^{N}$. It represents the Text Data of this $N$ time series. The timestamp descriptor $T_{i}$ converts the start/end timestamps of $\mathbf{s}_{i}$ into natural language phrases (e.g., “The time range of this sequence is from 03-Jan-2023 08:00 to 03-Jan-2023 12:00”) via $\operatorname{TimeStampDescriptor}(\cdot)$.
3.1.2 Statistical Descriptor
For statistical descriptor, the $S_{i}$ extracts distributional properties using $\operatorname{StatisticalDescriptor}(\cdot)$:
$\operatorname{StatisticalDescriptor}(\mathbf{s}_{i})=[\mu(\mathbf{s}_{i}),\sigma(\mathbf{s}_{i}),\nabla(\mathbf{s}_{i})]\in\mathbb{R}^{3}$ (2)
$S_{i}=\operatorname{StatisticalDescriptor}(\mathbf{s}_{i})$ (3)
where $\mu(\cdot)$, $\sigma(\cdot)$, and ($\nabla$($\cdot$)∇ ( ⋅ ) compute the mean, standard deviation, and the change in the sequence respectively, inspired by feature engineering in this work nie2022time .
After obtaining $T_{i}$ and $S_{i}$, we can splice them together:
$A_{i}=T_{i}\oplus S_{i}$ (4)
where $\oplus$ represents stitching $T_{i}$ and $S_{i}$ together and then we get $A_{i}$.
@@ 3.1.2 Statistical Descriptor
As shown in Fig. 3, this is the demonstration of the stitching operation. The $A_{i}$ is encoded through frozen LLM layers from pre-trained models dubey2024llama , with $\operatorname{SelectFinal}(\cdot)$ extracting the final-dimension representation $\mathbf{TE}_{i}\in\mathbb{R}^{D}$, where $D$ stands for LLM hidden dimension. This hybrid design ensures $\mathbf{TE}_{i}$ encodes both numerical regularity and contextual temporality, establishing cross-modal alignment while avoiding redundant LLM computations during training through offline precomputation jin2023time .
Figure 3: This prompt structure combines Timestamp Descriptor and Statistical Descriptor to systematically characterize time series data.
3.2 Adaptive Fusion Embedding Structure
To establish a synergistic interaction between numerical time series patterns and language-derived semantic context, we propose an adaptive fusion embedding structure that dynamically combines modality-specific embeddings through learnable gating. As shown in Fig. 2, this structure processes two complementary representations for each time series segment:
(1) Numerical Embedding ($\mathbf{SE}_{i}$): Captures intrinsic temporal dynamics through parameterized segment encoding:
$\mathbf{SE}_{i}=\operatorname{SegmentEmbedding}(\mathbf{s}_{i})\in\mathbb{R}^{D}$ (5)
where $\operatorname{SegmentEmbedding}(\cdot)$ implements learnable linear projections followed by GeLU activation hendrycks2016gaussian , transforming raw series segments into latent vectors aligned with LLM dimensions.
(2) Semantic Embedding ($\mathbf{TE}_{i}$): Encodes domain-specific statistical features and temporal context through frozen LLM processing of hybrid prompts:
$\mathbf{TE}_{i}=\operatorname{SelectFinal}\left(\operatorname{LLMLayers}\left(T_{i}\oplus S_{i}\right)\right)\in\mathbb{R}^{D}$ (6)
where $T_{i}$ denotes timestamp descriptors and $S_{i}$ contains statistical features as defined in Section 3.1.
The fusion process adaptively adjusts modality contributions through a learnable gating parameter $\theta\in\mathbb{R}$:
$\mathbf{E}_{i}=\underbrace{\sigma(\theta)}_{\alpha}\cdot\mathbf{SE}_{i}+\underbrace{(1-\sigma(\theta))}_{1-\alpha}\cdot\mathbf{TE}_{i}$ (7)
where $\sigma(\cdot)$ denotes the sigmoid function constraining $\alpha\in(0,1)$, the trainable $\theta$ automatically learns optimal fusion ratios from data, offering two key advantages such as dynamic adaptation to varying temporal regimes (e.g., prioritizing $\mathbf{SE}_{i}$ during stable periods while emphasizing $\mathbf{TE}_{i}$ when encountering anomalous patterns), and the bounded gating range prevents abrupt modality switching, with gradient flow governed by:
$\frac{\partial\mathbf{E}_{i}}{\partial\theta}=\sigma(\theta)(1-\sigma(\theta))(\mathbf{SE}_{i}-\mathbf{TE}_{i})$ (8)
This formulation ensures smooth gradient propagation to both modalities regardless of $\alpha$ values.
3.3 Dynamic Mixture-of-Experts Structure
To enhance the model’s capacity for capturing diverse temporal patterns, we design a dynamic mixture-of-experts structure that enables specialized feature learning through parameterized projection gates. The structure operates on contextualized embeddings $\{\hat{\mathbf{E}}_{i}\}\in\mathbb{R}^{B\times D}$ generated by LLM layers, where $B$ denotes the batch size and $D$ the hidden dimension.
The core innovation lies in training $K$ expert projections with sparsity constraints, where each expert $\mathbf{W}^{k}\in\mathbb{R}^{D\times D}$ learns distinct temporal dynamics. The gating structure computes modality selection weights through dimension-specific transformation: